# Limits of Probability Measures with General Coefficients

### v2

@@ 3.4. Remainders
This is clear by showing that the $N\rightarrow\infty$ limit of each of the finitely many terms of $\frac{1}{N^{k+\epsilon}}Q(c_{F}^{\nu}=a_{\nu}(N),\nu\in P^{+};N)$ is 00. ∎
Suppose we have a polynomial $R(c_{F}^{\nu},\nu\in P^{+};N)$ which is of order $N^{k}$. From the conditions of Theorem 3.2, we see that for each $\nu\in P^{+}$, there exists a constant $C>0$ such that $|c_{F_{N}}^{\nu}|\leq CN^{1-c}$ for sufficiently large $N\geq 1$. For all $\nu\in P^{+}$, let $a_{\nu}$ be the function such that $a_{\nu}(N)=c_{F_{N}}^{\nu}$ for $N\geq 1$. Afterwards, from Lemma 3.11, for $\epsilon>0$,
@@
We use (16) in the proof of Theorem 3.2 in Subsection 3.6 to show that such remainders $R$ have a $N\rightarrow\infty$ limit of 00. Particularly, (16) can be used with various results from Subsections 3.4 and 3.5.
@@
Suppose that in $l\in I_{N}(\lambda)$, the first $\lambda_{1}$ indices are $i_{1}$, the next $\lambda_{2}$ indices are $i_{2}$, and so forth, until the last $\lambda_{m}$ indices are $i_{m}$, where $1\leq i_{j}\leq N$ for $1\leq j\leq m$. From Proposition 3.10, for all $l\in I_{N}(\lambda)$, $\mathcal{D}^{\theta_{N}}_{l}(F(x_{1},\ldots,x_{N}))$ has order $N^{k(1-c)}$. Note that if $l$ has all $i_{j}$ distinct, $[1]\mathcal{D}^{\theta_{N}}_{l}(F(x_{1},\ldots,x_{N}))$ is equal, by symmetry, to $[1]\mathcal{D}^{\theta_{N}}_{\mathfrak{r}}(F(x_{1},\ldots,x_{N}))$. Consider the set $S$ of other $l$, where some of the $i_{j}$ are equal. Let
@@
From Corollary 3.6, $R^{\prime}$ is a polynomial in the $c_{F}^{\nu}$ which is homogeneous of degree $k$.
@@
$[1]\mathcal{D}^{\theta_{N}}_{\mathfrak{r}}(F(x_{1},\ldots,x_{N}))=\sum_{s\in H}r(s)_{k}+R$
@@
is order $N^{(k-m)(1-c)-1}$ from Definition 3.8. We consider the set $T$ of $s\in Q$ such that $r(s)_{k}$ equals $p$ multiplied by a nonzero integer. Following the proof of Proposition 3.10, let the coefficient of $p$ be $r$. We want to show that $|r|$ is $O(N^{(k-m)(1-c)-1})$. For $0\leq l\leq k-m$, suppose $T_{l}$ is the set of $s\in T$ with $l$ switches. For $s\in T_{l}$, we know that $|C(s)|\leq P_{l}\cdot N^{-lc}$ for a $P_{l}>0$ independent of $N$, see Proposition 3.9. Also, for $s\in T$, let $D(s)$ be the set of $x_{i}$ such that $i\notin\mathfrak{r}$ which are degree-altered by a term multiplication of $s$. Let $d=|D(s)|$.
If $s\in T_{l}$, $s$ has $m$ term multiplications, or $\alpha(s)$ operators. Also, there are $d$, $d\leq l$, $\beta(s)$ operators which are switches from $i_{j}$ to $i$ where $x_{i}\in D(s)$, such that each $x_{i}\in D(s)$ is in exactly one $\beta(s)$ operator. Note that the $\beta(s)$ operators exist by 3.10.1. The other $k-m-d$ operators are $\gamma(s)$ operators, and are derivatives or switches. Note that the number of switches in the $\gamma(s)$ operators is $l-d$, and each of these switch has $N-1$ possibilities.
@@ 3.5. Sequences involving changes
Next, we consider the operators $\mathcal{Q}_{i}^{N}$ from Section 2, which involve changes. Recall that the operators are defined in (6). Suppose that for $N\geq i$,
@@
is considered. In Theorem 3.17, we show that we can replace each $\mathcal{D}^{\theta_{N}}_{i_{j}}$ with $N^{1-c}\mathcal{Q}_{i_{j}}^{N}\left(\frac{\partial_{i_{j}}F}{N^{1-c}}\right)$. Afterwards, in Theorem 3.20, we use the $\mathcal{R}_{i,j}^{k}$ operators to evaluate the resulting expression. Similarly as before, for $\mathfrak{r}=\{i_{j}\}_{1\leq j\leq k}$ and $N\geq\max(\mathfrak{r})$, we let
@@
$[1]\mathcal{Q}_{\mathfrak{r}}^{N}(F(x_{1},\ldots,x_{N}))=\sum_{s\in T_{N,\theta_{N}}^{2}(\mathfrak{r})}r(s)_{k}.$
The following analogue of Corollary 3.6 holds.
@@
For a sequence $\mathfrak{r}=\{i_{j}\}_{1\leq j\leq k}$ of positive integers and $N\geq\max(\mathfrak{r})$, $[1]\mathcal{Q}^{N}_{\mathfrak{r}}(F(x_{1},\ldots,x_{N}))$ is a polynomial in the $c_{F}^{\nu}$ for $\nu\in P^{+}$ which is homogeneous of degree $k$, where $c_{F}^{\nu}$ has degree $|\nu|$ for $\nu\in P^{+}$.
@@
This follows from summing the applications of Lemma 3.5 to sequences $s$ with indices $\mathfrak{r}$ and factor $\theta_{N}$ that only contain term multiplications and changes. ∎
@@
Let $\mathfrak{r}=\{i_{j}\}_{1\leq j\leq k}$ be a sequence of positive integers. For $N\geq\max(\mathfrak{r})$, $[1]\mathcal{Q}^{N}_{\mathfrak{r}}(F(x_{1},\ldots,x_{N}))$ has order $N^{k(1-c)}$.
@@
The same proof as for Proposition 3.10 can be used, but with no derivatives and changes instead of switches. Particularly, if a switch is from $i$ to $j$, a change from $i$ to $j$ is used instead. ∎
@@
$[1]\mathcal{Q}_{\mathfrak{r}}^{N}(F(x_{1},\ldots,x_{N}))=\sum_{s\in H}r(s)_{k}+R$
@@
The same proof as for Lemma 3.13 can be used, but with no derivatives and changes instead of switches. Particularly, if a switch is from $i$ to $j$, a change from $i$ to $j$ is used instead. ∎
@@
(17) $[1]\mathcal{D}^{\theta_{N}}_{\mathfrak{r}}(F(x_{1},\ldots,x_{N}))=[1]\mathcal{Q}_{\mathfrak{r}}^{N}(F(x_{1},\ldots,x_{N}))+R$
@@
Suppose that $H$ is the set of $s$ in $T_{N,\theta_{N}}^{1}(\mathfrak{r})$ satisfying the Lemma 3.13 conditions. From the Lemma,
$[1]\mathcal{D}^{\theta_{N}}_{\mathfrak{r}}(F(x_{1},\ldots,x_{N}))=\sum_{s\in H}r(s)_{k}+R_{1},$
where $R_{1}$ is a homogeneous polynomial in the $c_{F}^{\nu}$ with degree $k$ and order $N^{k(1-c)-1}$. Consider $H_{1}=H\cap T_{N,\theta_{N}}^{3}(\mathfrak{r})$. We see that $H_{1}$ is the set of $s\in T_{N,\theta_{N}}^{1}(\mathfrak{r})$ satisfying the conditions of Lemma 3.13 with only switches and term multiplications, and
@@
$[1]\mathcal{D}^{\theta_{N}}_{\mathfrak{r}}(F(x_{1},\ldots,x_{N}))=\sum_{s\in H_{1}}r(s)_{k}+R_{1}+R_{2}.$
Also, let $H_{2}$ be the set of $s\in T_{N,\theta_{N}}^{2}(\mathfrak{r})$ satisfying the Lemma 3.16 conditions. By the lemma,
$[1]\mathcal{Q}_{\mathfrak{r}}^{N}(F(x_{1},\ldots,x_{N}))=\sum_{s\in H_{2}}r(s)_{k}+R_{3},$
@@
Suppose $s\in T_{l}$. Note that the derivatives of $s$ only appear in $\gamma(s)$ operators. Suppose $|D(s)|=d\leq l$. We know that in the $\gamma(s)$ operators, we have $l-d$ switches and $k-m-l$ derivatives. Since the $\gamma(s)$ operators must have at least $1$ derivative, we see that $0\leq l<k-m$. Particularly, $|T_{l}|$ is $O(N^{l})$ for $0\leq l<k-m$ and $|T_{l}|=0$ for $l\geq k-m$, after using the method of Lemma 3.13. Since $|C(s)|$ is $O(N^{-lc})$ for $s\in T_{l}$ by Proposition 3.9,
@@
Suppose $q$ is a term of $r(s)_{j-1}$. We look at $s_{j}(q)$, which is a sum of terms or 00. Suppose $q^{\prime}$ is a term of $s_{j}(q)$ and the degree of $x_{i}$ in $q^{\prime}$ is at least $1$. For $j<j^{\prime}\leq k$, $s_{j^{\prime}}$ is the switch from $i_{j^{\prime}}$ to $i^{\prime}\not=i$ or a term multiplication. In $r(s)_{k}$, the degree of $x_{i}$ must be 00, but to decrease the degree of $x_{i}$ of a term, some $s_{j^{\prime}}$ for $j^{\prime}>j$ will convert the term to 00. Therefore, $q^{\prime}$ will not contribute to $r(s)_{k}$. So, the only term of $s_{j}(q)$ which will contribute to $r(s)_{k}$ will be $q^{\prime}$ with degree of $x_{i}$ equal to 00, and since $s_{j}$ is the switch from $i_{j}$ to $i$, in $s_{j}(q)$, there will be at most one such $q^{\prime}$. With this, we can set $s_{j}(q)$ as $q^{\prime}$ if such $q^{\prime}$ exists and $s_{j}(q)=0$ if not.
@@
Next, suppose the degree of $x_{i}$ is 00 and $x_{i_{j}}$ is at least $1$ in $q$. In this case, $q^{\prime}$ exists and will be $\theta_{N}d_{i_{j}}(q)$. But, since the degree of $x_{i_{j}}$ is at least $1$,
@@
Consider when the degree of $x_{i_{j}}$ in $q$ is 00. If the degree of $x_{i}$ is at least $1$ in $q$, then $q^{\prime}$ exists and is equal to $-\theta_{N}C_{i_{j},i}(q)$. Otherwise, if the degree of $x_{i}$ is 00, $s_{j}(q)=0=-\theta_{N}C_{i_{j},i}(q)$. However, since the degree of $x_{i_{j}}$ in $q$ is 00,
@@
Note that the proof of the previous claim has similarities with the argument that appears in step 3 of the proof of Proposition 5.5 of [matrix] that justifies replacing a switch from $i_{j}$ to $i$ with $d_{i}$. In contrast, we replace a switch from $i_{j}$ to $i$ with a change from $i_{j}$ to $i$. The reason for this is that there may be terms that contain $x_{i}$ prior to the application of the switch from $i_{j}$ to $i$ due to the nonzero limits of coefficients of terms with multiple indices, which requires the usage of the $C_{i_{j},i}$ operators. To address this difference, we analyze the $\mathcal{Q}_{\mathfrak{r}}^{N}$ operators in Section 2.
Since $c:H_{1}\rightarrow H_{2}$ is a bijection, with 3.17.2, $\sum_{s\in H_{1}}r(s)_{k}=\sum_{s\in H_{2}}r(s)_{k}$. Therefore,
$[1]\mathcal{D}^{\theta_{N}}_{\mathfrak{r}}(F(x_{1},\ldots,x_{N}))=[1]\mathcal{Q}_{\mathfrak{r}}^{N}(F(x_{1},\ldots,x_{N}))+R_{1}+R_{2}-R_{3}.$
@@
The main difference between the regime $c<1$ that we consider and the regime $c=1$ arises from the order of the derivative operator. As implied by 3.17.1, when $c<1$, we can disregard the derivative operators entirely because they are of lower order.
@@
Suppose that $\lambda$ is a partition with $\ell(\lambda)=m$ and $|\lambda|=k$. Suppose $N\geq m$. For $1\leq i\leq N$, let $\vec{X}_{i}=(x_{1},\ldots,x_{i},0,\ldots,0)$ be $\vec{x}$ with $x_{j}$ replaced by 00 for $i+1\leq j\leq N$. Then, where $\mathfrak{r}$ is the sequence of indices from Proposition 3.12,
@@
$[1]N^{k(1-c)}\prod_{i=1}^{m}\mathcal{Q}_{i}^{N}\left(\frac{\partial_{i}F}{N^{1-c}}\right)^{\lambda_{i}}(1)=[1]N^{k(1-c)}\prod_{i=1}^{m}\mathcal{Q}_{N-i+1}^{N}\left(\frac{\partial_{N-i+1}F}{N^{1-c}}\right)^{\lambda_{i}}(1).$
Here, $\mathfrak{r}=\{i_{j}\}_{1\leq j\leq k}$ consists of $k$ indices, the first $\lambda_{1}$ being $N$, the next $\lambda_{2}$ being $N-1$, and so forth, until the last $\lambda_{m}$ are $N-m+1$. Let $T$ be the set of sequences $s$ with indices $\mathfrak{r}$ over $x_{1},\ldots,x_{N}$ consisting of term multiplications and changes such that:
  • For $11\leq j\leq m-$ 1, after $\lambda_{1}+\cdots+\lambda_{j}$ operators, the next $\lambda_{j+1}$ operators cannot have changes from $N-j$ to $i$ and term multiplications which degree-alter $x_{i}$, $N-j+1\leq i\leq N$.
  • $r(s)_{k}$ is 00 or has degree 00 in the $x_{i}$.
@@
To show this, we prove that $r_{d}(s)_{k}=r_{d-1}(s)_{k}$, $1\leq d\leq m$. Observe that $r_{d}(s)_{j}=r_{d-1}(s)_{j}$ for $11\leq j\leq\lambda_{1}+\cdots+\lambda_{d}-$ 1 and
@@
where each term of $r_{d-1}(s)_{\lambda_{1}+\cdots+\lambda_{d}}-r_{d-1}(s)_{\lambda_{1}+\cdots+\lambda_{d}}|_{x_{N-d+1}=0}$ contains $x_{N-d+1}$. But, if $\lambda_{1}+\cdots+\lambda_{d}+1\leq j\leq k$, $i_{j}\not=N-d+1$ since the indices of $s$ are $\mathfrak{r}$ and $s_{j}$ is not a change from $i_{j}$ to $N-d+1$ since $s\in T$. Hence, $s_{j}$ cannot remove $x_{N-d+1}$ from a term that contains $x_{N-d+1}$ without causing the term to become 00. Then, all nonzero terms of $r_{d-1}(s)_{k}-r_{d}(s)_{k}$ contain $x_{N-d+1}$. Since $s\in T$, $r_{0}(s)_{k}=r(s)_{k}$ has degree 00 in the $x_{i}$, so $r_{d-1}(s)_{k}$ and $r_{d}(s)_{k}$, and hence $r_{d-1}(s)_{k}-r_{d}(s)_{k}$, have degree 00 in the $x_{i}$ from 3.20.1. But, all nonzero terms of $r_{d-1}(s)_{k}-r_{d}(s)_{k}$ contain $x_{N-d+1}$, and therefore, $r_{d-1}(s)_{k}-r_{d}(s)_{k}=0$. ∎
@@
$[1]\prod_{i=1}^{m}\left((N-i+1)^{\lambda_{i}(1-c)}\mathcal{R}_{N-i+1,N-i+1}^{\lambda_{i}}\left(\frac{\partial_{N-i+1}F(\vec{X}_{N-i+1})}{(N-i+1)^{1-c}}\right)\right)(1)=\sum_{s\in T\cap T_{N,\theta_{N}}^{2}(\mathfrak{r})}r(s)_{k}.$
@@
since $r_{m}(s)_{k}=r(s)_{k}$ for $s\in T$ by 3.20.2 and $r(s)_{k}\not=0$ only if $s\in T\cap T_{N,\theta_{N}}^{2}(\mathfrak{r})$. ∎
Suppose that $T^{\prime}$ is the set of $s\in T_{N,\theta_{N}}^{2}(\mathfrak{r})$ satisfying the Lemma 3.16 conditions. From the lemma,
$[1]N^{k(1-c)}\prod_{i=1}^{m}\mathcal{Q}_{N-i+1}^{N}\left(\frac{\partial_{N-i+1}F}{N^{1-c}}\right)^{\lambda_{i}}(1)=\sum_{s\in T^{\prime}}r(s)_{k}+R_{1}$

### v3

@@ 3.4. Remainders
This is clear by showing that the $N\rightarrow\infty$ limit of each of the finitely many terms of $\frac{1}{N^{k+\epsilon}}Q(c_{F}^{\nu}=a_{\nu}(N),\nu\in P^{+};N)$ is 0. ∎
Suppose we have a polynomial $R(c_{F}^{\nu},\nu\in P^{+};N)$ which is of order $N^{k}$. From the conditions of Theorem˜3.2, we see that for each $\nu\in P^{+}$, there exists a constant $C>0$ such that $|c_{F_{N}}^{\nu}|\leq CN^{1-c}$ for sufficiently large $N\geq 1$. For all $\nu\in P^{+}$, let $a_{\nu}$ be the function such that $a_{\nu}(N)=c_{F_{N}}^{\nu}$ for $N\geq 1$. Afterwards, from Lemma˜3.11, for $\epsilon>0$,
@@
We use (15) in the proof of Theorem˜3.2 in Subsection 3.6 to show that such remainders $R$ have a $N\rightarrow\infty$ limit of 0. Particularly, (15) can be used with various results from Subsections 3.4 and 3.5.
@@
Suppose that in $l\in I_{N}(\lambda)$, the first $\lambda_{1}$ indices are $i_{1}$, the next $\lambda_{2}$ indices are $i_{2}$, and so forth, until the last $\lambda_{m}$ indices are $i_{m}$, where $1\leq i_{j}\leq N$ for $1\leq j\leq m$. From Proposition˜3.10, for all $l\in I_{N}(\lambda)$, $\mathcal{D}^{\theta_{N}}_{l}(F(x_{1},\ldots,x_{N}))$ has order $N^{k(1-c)}$. Note that if $l$ has all $i_{j}$ distinct, $1]\mathcal{D}^{\theta_{N}}_{l}(F(x_{1},\ldots,x_{N}))$ is equal, by symmetry, to $1]\mathcal{D}^{\theta_{N}}_{\mathfrak{r}}(F(x_{1},\ldots,x_{N}))$. Consider the set $S$ of other $l$, where some of the $i_{j}$ are equal. Let
@@
From Corollary˜3.6, $R^{\prime}$ is a polynomial in the $c_{F}^{\nu}$ which is homogeneous of degree $k$.
@@
$1]\mathcal{D}^{\theta_{N}}_{\mathfrak{r}}(F(x_{1},\ldots,x_{N}))=\sum_{s\in H}r(s)_{k}+R$
@@
is order $N^{(k-m)(1-c)-1}$ from Definition˜3.8. We consider the set $T$ of $s\in Q$ such that $r(s)_{k}$ equals $p$ multiplied by a nonzero integer. Following the proof of Proposition˜3.10, let the coefficient of $p$ be $r$. We want to show that $|r|$ is $O(N^{(k-m)(1-c)-1})$. For $0\leq l\leq k-m$, suppose $T_{l}$ is the set of $s\in T$ with $l$ switches. For $s\in T_{l}$, we know that $|C(s)|\leq P_{l}\cdot N^{-lc}$ for a $P_{l}>0$ independent of $N$, see Proposition˜3.9. Also, for $s\in T$, let $D(s)$ be the set of $x_{i}$ such that $i\notin\mathfrak{r}$ which are degree-altered by a term multiplication of $s$. Let $d=|D(s)|$.
If $\in T_{l}, ss$ has $m$ term multiplications, or $\alpha(s)$ operators. Also, there are $d$, $d\leq l$, $\beta(s)$ operators which are switches from $i_{j}$ to $i$ where $x_{i}\in D(s)$, such that each $x_{i}\in D(s)$ is in exactly one $\beta(s)$ operator. Note that the $\beta(s)$ operators exist by ˜3.10.1. The other kmdk-m-d operators are $\gamma(s)$ operators, and are derivatives or switches. Note that the number of switches in the $\gamma(s)$ operators is ldl-d, and each of these switch has N1N-1 possibilities.
@@ 3.5. Sequences involving changes
Next, we consider the operators $\mathcal{Q}_{i}^{N}$ from Section˜2, which involve changes. Recall that the operators are defined in (5). Suppose that for $N\geq i$,
@@
is considered. In Theorem˜3.17, we show that we can replace each $\mathcal{D}^{\theta_{N}}_{i_{j}}$ with $N^{1-c}\mathcal{Q}_{i_{j}}^{N}\left(\frac{\partial_{i_{j}}F}{N^{1-c}}\right)$. Afterwards, in Theorem˜3.20, we use the $\mathcal{R}_{i,j}^{k}$ operators to evaluate the resulting expression. Similarly as before, for $\mathfrak{r}=\{i_{j}\}_{1\leq j\leq k}$ and $N\geq\max(\mathfrak{r})$, we let
@@
$1]\mathcal{Q}_{\mathfrak{r}}^{N}(F(x_{1},\ldots,x_{N}))=\sum_{s\in T_{N,\theta_{N}}^{2}(\mathfrak{r})}r(s)_{k}.$
The following analogue of Corollary˜3.6 holds.
@@
For a sequence $\mathfrak{r}=\{i_{j}\}_{1\leq j\leq k}$ of positive integers and $N\geq\max(\mathfrak{r})$, $1]\mathcal{Q}^{N}_{\mathfrak{r}}(F(x_{1},\ldots,x_{N}))$ is a polynomial in the $c_{F}^{\nu}$ for $\nu\in P^{+}$ which is homogeneous of degree $k$, where $c_{F}^{\nu}$ has degree $|\nu|$ for $\nu\in P^{+}$.
@@
This follows from summing the applications of Lemma˜3.5 to sequences $s$ with indices $\mathfrak{r}$ and factor $\theta_{N}$ that only contain term multiplications and changes. ∎
@@
Let $\mathfrak{r}=\{i_{j}\}_{1\leq j\leq k}$ be a sequence of positive integers. For $N\geq\max(\mathfrak{r})$, $1]\mathcal{Q}^{N}_{\mathfrak{r}}(F(x_{1},\ldots,x_{N}))$ has order $N^{k(1-c)}$.
@@
The same proof as for Proposition˜3.10 can be used, but with no derivatives and changes instead of switches. Particularly, if a switch is from $i$ to $j$, a change from $i$ to $j$ is used instead. ∎
@@
$1]\mathcal{Q}_{\mathfrak{r}}^{N}(F(x_{1},\ldots,x_{N}))=\sum_{s\in H}r(s)_{k}+R$
@@
The same proof as for Lemma˜3.13 can be used, but with no derivatives and changes instead of switches. Particularly, if a switch is from $i$ to $j$, a change from $i$ to $j$ is used instead. ∎
@@
(16) $1]\mathcal{D}^{\theta_{N}}_{\mathfrak{r}}(F(x_{1},\ldots,x_{N}))=[1]\mathcal{Q}_{\mathfrak{r}}^{N}(F(x_{1},\ldots,x_{N}))+R$
@@
Suppose that $H$ is the set of $s$ in $T_{N,\theta_{N}}^{1}(\mathfrak{r})$ satisfying the Lemma˜3.13 conditions. From the Lemma,
$1]\mathcal{D}^{\theta_{N}}_{\mathfrak{r}}(F(x_{1},\ldots,x_{N}))=\sum_{s\in H}r(s)_{k}+R_{1},$
where $R_{1}$ is a homogeneous polynomial in the $c_{F}^{\nu}$ with degree $k$ and order $N^{k(1-c)-1}$. Consider $H_{1}=H\cap T_{N,\theta_{N}}^{3}(\mathfrak{r})$. We see that $H_{1}$ is the set of $s\in T_{N,\theta_{N}}^{1}(\mathfrak{r})$ satisfying the conditions of Lemma˜3.13 with only switches and term multiplications, and
@@
$1]\mathcal{D}^{\theta_{N}}_{\mathfrak{r}}(F(x_{1},\ldots,x_{N}))=\sum_{s\in H_{1}}r(s)_{k}+R_{1}+R_{2}.$
Also, let $H_{2}$ be the set of $s\in T_{N,\theta_{N}}^{2}(\mathfrak{r})$ satisfying the Lemma˜3.16 conditions. By the lemma,
$1]\mathcal{Q}_{\mathfrak{r}}^{N}(F(x_{1},\ldots,x_{N}))=\sum_{s\in H_{2}}r(s)_{k}+R_{3},$
@@
Suppose $s\in T_{l}$. Note that the derivatives of $s$ only appear in $\gamma(s)$ operators. Suppose $|D(s)|=d\leq l$. We know that in the $\gamma(s)$ operators, we have ldl-d switches and kmlk-m-l derivatives. Since the $\gamma(s)$ operators must have at least $1$ derivative, we see that $0\leq l<k-m$. Particularly, $|T_{l}|$ is $O(N^{l})$ for $0\leq l<k-m$ and $|T_{l}|=0$ for $l\geq k-m$, after using the method of Lemma˜3.13. Since $|C(s)|$ is $O(N^{-lc})$ for $s\in T_{l}$ by Proposition˜3.9,
@@
Suppose $q$ is a term of $r(s)_{j-1}$. We look at $s_{j}(q)$, which is a sum of terms or 0. Suppose $q^{\prime}$ is a term of $s_{j}(q)$ and the degree of $x_{i}$ in $q^{\prime}$ is at least $1$. For $j<j^{\prime}\leq k$, $s_{j^{\prime}}$ is the switch from $i_{j^{\prime}}$ to $i^{\prime}\not=i$ or a term multiplication. In $r(s)_{k}$, the degree of $x_{i}$ must be 0, but to decrease the degree of $x_{i}$ of a term, some $s_{j^{\prime}}$ for $j^{\prime}>j$ will convert the term to 0. Therefore, $q^{\prime}$ will not contribute to $r(s)_{k}$. So, the only term of $s_{j}(q)$ which will contribute to $r(s)_{k}$ will be $q^{\prime}$ with degree of $x_{i}$ equal to 0, and since $s_{j}$ is the switch from $i_{j}$ to $i$, in $s_{j}(q)$, there will be at most one such $q^{\prime}$. With this, we can set $s_{j}(q)$ as $q^{\prime}$ if such $q^{\prime}$ exists and $s_{j}(q)=0$ if not.
@@
Next, suppose the degree of $x_{i}$ is 0 and $x_{i_{j}}$ is at least $1$ in $q$. In this case, $q^{\prime}$ exists and will be $\theta_{N}d_{i_{j}}(q)$. But, since the degree of $x_{i_{j}}$ is at least $1$,
@@
Consider when the degree of $x_{i_{j}}$ in $q$ is 0. If the degree of $x_{i}$ is at least $1$ in $q$, then $q^{\prime}$ exists and is equal to $\theta_{N}C_{i_{j},i}(q)$. Otherwise, if the degree of $x_{i}$ is 0, $s_{j}(q)=0=-\theta_{N}C_{i_{j},i}(q)$. However, since the degree of $x_{i_{j}}$ in $q$ is 0,
@@
Note that the proof of the previous claim has similarities with the argument that appears in step 3 of the proof of Proposition 5.5 of [matrix] that justifies replacing a switch from $i_{j}$ to $i$ with $d_{i}$. In contrast, we replace a switch from $i_{j}$ to $i$ with a change from $i_{j}$ to $i$. The reason for this is that there may be terms that contain $x_{i}$ prior to the application of the switch from $i_{j}$ to $i$ due to the nonzero limits of coefficients of terms with multiple indices, which requires the usage of the $C_{i_{j},i}$ operators. To address this difference, we analyze the $\mathcal{Q}_{\mathfrak{r}}^{N}$ operators in Section˜2.
Since $c:H_{1}\rightarrow H_{2}$ is a bijection, with ˜3.17.2, $\sum_{s\in H_{1}}r(s)_{k}=\sum_{s\in H_{2}}r(s)_{k}$. Therefore,
$1]\mathcal{D}^{\theta_{N}}_{\mathfrak{r}}(F(x_{1},\ldots,x_{N}))=[1]\mathcal{Q}_{\mathfrak{r}}^{N}(F(x_{1},\ldots,x_{N}))+R_{1}+R_{2}-R_{3}.$
@@
The main difference between the regime $c<1$ that we consider and the regime $c=1$ arises from the order of the derivative operator. As implied by ˜3.17.1, when $c<1$, we can disregard the derivative operators entirely because they are of lower order.
@@
Suppose that $\lambda$ is a partition with $\ell(\lambda)=m$ and $|\lambda|=k$. Suppose $N\geq m$. For $1\leq i\leq N$, let $\vec{X}_{i}=(x_{1},\ldots,x_{i},0,\ldots,0)$ be $\vec{x}$ with $x_{j}$ replaced by 0 for $i+1\leq j\leq N$. Then, where $\mathfrak{r}$ is the sequence of indices from Proposition˜3.12,
@@
$1]N^{k(1-c)}\prod_{i=1}^{m}\mathcal{Q}_{i}^{N}\left(\frac{\partial_{i}F}{N^{1-c}}\right)^{\lambda_{i}}(1)=[1]N^{k(1-c)}\prod_{i=1}^{m}\mathcal{Q}_{N-i+1}^{N}\left(\frac{\partial_{N-i+1}F}{N^{1-c}}\right)^{\lambda_{i}}(1).$
Here, $\mathfrak{r}=\{i_{j}\}_{1\leq j\leq k}$ consists of $k$ indices, the first $\lambda_{1}$ being $N$, the next $\lambda_{2}$ being N1N-1, and so forth, until the last $\lambda_{m}$ are Nm+1N-m+1. Let $T$ be the set of sequences $s$ with indices $\mathfrak{r}$ over $x_{1},\ldots,x_{N}$ consisting of term multiplications and changes such that:
  • For $1\leq j\leq m-1$, after $\lambda_{1}+\cdots+\lambda_{j}$ operators, the next $\lambda_{j+1}$ operators cannot have changes from NjN-j to $i$ and term multiplications which degree-alter $x_{i}$, $N-j+1\leq i\leq N$.
  • $r(s)_{k}$ is 0 or has degree 0 in the $x_{i}$.
@@
To show this, we prove that $r_{d}(s)_{k}=r_{d-1}(s)_{k}$, $1\leq d\leq m$. Observe that $r_{d}(s)_{j}=r_{d-1}(s)_{j}$ for $1\leq j\leq\lambda_{1}+\cdots+\lambda_{d}-1$ and
@@
where each term of $r_{d-1}(s)_{\lambda_{1}+\cdots+\lambda_{d}}-r_{d-1}(s)_{\lambda_{1}+\cdots+\lambda_{d}}|_{x_{N-d+1}=0}$ contains $x_{N-d+1}$. But, if $\lambda_{1}+\cdots+\lambda_{d}+1\leq j\leq k$, $i_{j}\not=N-d+1$ since the indices of $s$ are $\mathfrak{r}$ and $s_{j}$ is not a change from $i_{j}$ to Nd+1N-d+1 since $s\in T$. Hence, $s_{j}$ cannot remove $x_{N-d+1}$ from a term that contains $x_{N-d+1}$ without causing the term to become 0. Then, all nonzero terms of $r_{d-1}(s)_{k}-r_{d}(s)_{k}$ contain $x_{N-d+1}$. Since $s\in T$, $r_{0}(s)_{k}=r(s)_{k}$ has degree 0 in the $x_{i}$, so $r_{d-1}(s)_{k}$ and $r_{d}(s)_{k}$, and hence $r_{d-1}(s)_{k}-r_{d}(s)_{k}$, have degree 0 in the $x_{i}$ from ˜3.20.1. But, all nonzero terms of $r_{d-1}(s)_{k}-r_{d}(s)_{k}$ contain $x_{N-d+1}$, and therefore, $r_{d-1}(s)_{k}-r_{d}(s)_{k}=0$. ∎
@@
$1]\prod_{i=1}^{m}\left((N-i+1)^{\lambda_{i}(1-c)}\mathcal{R}_{N-i+1,N-i+1}^{\lambda_{i}}\left(\frac{\partial_{N-i+1}F(\vec{X}_{N-i+1})}{(N-i+1)^{1-c}}\right)\right)(1)=\sum_{s\in T\cap T_{N,\theta_{N}}^{2}(\mathfrak{r})}r(s)_{k}.$
@@
since $r_{m}(s)_{k}=r(s)_{k}$ for $s\in T$ by ˜3.20.2 and $r(s)_{k}\not=0$ only if $s\in T\cap T_{N,\theta_{N}}^{2}(\mathfrak{r})$. ∎
Suppose that $T^{\prime}$ is the set of $s\in T_{N,\theta_{N}}^{2}(\mathfrak{r})$ satisfying the Lemma˜3.16 conditions. From the lemma,
$1]N^{k(1-c)}\prod_{i=1}^{m}\mathcal{Q}_{N-i+1}^{N}\left(\frac{\partial_{N-i+1}F}{N^{1-c}}\right)^{\lambda_{i}}(1)=\sum_{s\in T^{\prime}}r(s)_{k}+R_{1}$